# Automated Proof Generation for Rust Code via Self-Evolution

### v1

@@ 3.1 Step 1: Generating Verus-compatible Code
To address this compatibility issue, we employ GPT-4o as a code translator, effectively substituting Verus-incompatible Rust code snippets with Verus-compatible alternatives, such as converting an iterator-based implemenation into a while-loop based implementation. Specifically, we prompt GPT-4o with all the unsupported features sourced from Verus official documentation. All the converted codes that can pass the grammar check of Verus compiler are collected for next steps. This step is performed only once in our self-evolving framework.
@@ 3.2 Step 2: Self-evolving Specification Synthesis
In this step, LLMs are tasked with generating preconditions and postconditions for a Rust function based on the function’s implementation and docstring. Different from prior work in spec-generation (Flanagan & Leino, 2001; Ma et al., 2024), SAFE needs synthesized specification as input for its self-evolving proof-synthesis framework, which raises unique requirements on its spec-evaluation criteria (i.e., which specification to keep or discard) and the mechanism to evaluate the criteria.
@@
In terms of criteria, generally speaking, a perfect specification $S$ should be correct (i.e., any correct implementation should be accepted by $S$) and complete (i.e., any incorrect code implementation should be rejected by $S$). For example, the spec in the left column below is correct but incomplete for the binary search function, as an incorrect implementation that always returns 0 would be accepted by this spec; the spec in the right column below is incorrect, as a correct implementation may fail this spec when the input array has multiple elements matching the search key. In SAFE, getting perfect specification is not the goal. SAFE should discard incorrect specifications, which can never be proved no matter what proof annotations are used and hence will waste training cycles in the next step. However, SAFE is fine with incomplete specifications, as the usage of spec in SAFE is to stimulate proof synthesis not to judge code correctness. In fact, SAFE needs incomplete specifications, as complete specifications are likely too challenging to prove at the early stage of the proof-synthesis model’s evolution. Of course, specifications that are too incomplete should be discarded, because trivial specifications can be proved without any proof annotations and hence offer no usage for proof-synthesis training (e.g., 0 <= r for the binary search function).
1#A correct, but incomplete specification.
20 <= r,
3r < v.len(),
1#An incorrect specification.
2r > 0 ==> k > v[r - 1],
3r < v.len() - 1 ==> k < v[r + 1],
@@
In terms of the mechanism to evaluate the above criteria, much previous work relies on running many test cases, formal verifying the consistency between spec and code, or user inspection. None of these mechanisms suit SAFE: relying on user inspection or running the program through many test cases would severely slow down the self-evolving framework; formal verification also does not work when we do not have a capable model to automatically synthesize proof.
@@
With all these constraints in mind, SAFE chooses to leverage a recently proposed technique (Lahiri, 2024) to symbolically and quantitatively measure the Correctness and Completeness of specification. Specifically, we use Verus to symbolically evaluate what percentage of test cases $\mathbf{T}$ provided by the original dataset can pass a given specification $S$ (i.e., the Correctness score), and what percentage of test cases $\mathbf{T^{\prime}}$ mutated from $\mathbf{T}$ can be rejected by $S$ (i.e., the Completeness score). Since Verus can symbolically conduct this evaluation without checking code implementation, this process can be done very quickly; since this technique offers quantitative scores, we can set a threshold that fits the unique need of SAFE. Specifically, we keep all the specifications that have higher than 80% of Correctness score and higher than 60% of Completeness score. This way, SAFE can obtain many imperfect but useful specifications for its proof synthesis. We do not set the Correctness threshold to be 100%, because we also use GPT to generate some test cases when few test cases are provided by the dataset, and we cannot guarantee perfect correctness of these test cases.
@@
We keep up to three synthesized specifications for each function, which not only enriches the spec-training dataset but also increases the chance of synthesizing valid proof later. Our self-evolving process for specification stops when there is no significant increase in total number of accepted function–spec pairs, at which time all these pairs move on to become the input for the next step.
@@ 3.3 Step 3: Self-evolving Proof Synthesis
Now that we have tens of thousands of function-specification pairs, we can start the self-evolving procedure of proof synthesis. Comparing with spec-synthesis, deciding whether to accept or reject a synthesized item is much more straightforward — only proof that Verus can use to verify the Rust function satisfies its specification is accepted; if any verification errors are raised, the proof is discarded [^1] If a Rust function is in fact buggy or a synthesized specification is incorrect, no proof will be accepted.. However, since proof synthesis is much more difficult than specification synthesis, we need to pay special attention to bootstrap the self-evolving procedure and to keep it moving forward.
For bootstrapping, we realized that a simple prompt to GPT-4o, like “You are an expert in Verus. Your mission is to write proof annotations for the following function”, would not work. It is very difficult for GPT-4o to synthesize all the needed loop invariants without any incorrect ones in between. More complicated proof annotations like proof blocks and lemma functions are even more out of reach. Therefore, we write a detailed prompt for GPT-4o, explaining the principles and tricks of writing Verus proof. Designing these prompts took three Verus experts about 40 working hours. By doing so, we finally manage to make GPT-4o synthesize proof annotations that allow Verus to prove more than one thousand Rust functions at the cost of one whole month of non-stop GPT-4o invocation. Fortunately, this is sufficient for bootstrap. For the remainder of the self-evolving procedure, GPT-4o is not used any more. Instead, we use an open-source LLM much smaller than GPT-4o to efficiently synthesize data and improve its capability in a self-evolving manner as in Algorithm 1.
@@ 4.1.1 Dataset
As an early exploration of automated formal verification, in this paper, we focus solely on the Rust code of algorithm types at the function level, as our source of data. We employ the MBPP dataset (Austin et al., 2021) (only training split for data synthesis) and the CodeNet dataset (Puri et al., 2021) as our data sources. These two datasets contain small programs written in different programming languages, each of which is an intended solution to a coding problem described in natural language and is associated with three test cases on average. We translate the Python programs in MBPP and extract the Rust programs in CodeNet. In total, we collect 45,395 Rust single-function programs. Of these, 21,398 have been successfully transformed into Rust code that is compatible with Verus by SAFE. We conduct the specification synthesis first, after two rounds of self-evolution, we obtain 19,017 high-quality specifications. Then we run three rounds of self-evolving proof synthesis, ending with 9,706 verified programs, and 10,486 self-debugging data pairs. To the best of our knowledge, this represents the most extensive synthetic dataset created for Verus code to date.
4.1.2 Benchmark and Metrics
We assess the model’s proficiency in generating proofs in two benchmarks, the human-written benchmark VerusBench and the synthetic benchmark CodeNet-Test.
VerusBench is a human-written benchmark dataset with human-written Rust code and corresponding Verus specifications. It contains 139 code files in total. 23 of them are algorithmic programs from Verus tutorials. 38 of them come from the dataset of SV-COMP-2021 (Beyer, 2021) (in short, SV) – a contest focused on program verification for C and Java languages; we utilize 38 Verus-translated tasks provided by (Yao et al., 2023). The remaining 78 tasks come from MBPP-DFY-153 (Misu et al., 2024), a Dafny version of the MBPP test set. We have translated 78 of them that are compatible with Verus.
@@ 4.1.2 Benchmark and Metrics
CodeNet-Test is an expansive benchmark, 10x larger than VerusBench, crafted by LLM and encompassing 1,435 diverse tasks. Given the substantial human effort required to write Verus code and specifications, it’s impractical to create a large test suite by human. To comprehensively measure the model capability of generating proofs, we split a subset of CodeNet for testing, ensuring it was not utilized in our data synthesis process. We leverage our specification generator to craft specifications for this subset and apply the same scoring mechanism to preserve reasonably well specifications.
@@
We measure the capability of proof generation by Accuracy@K. Specifically, we use our model to sample K proofs during proof generation, and for each incorrect sample, we sample K proofs generated during self-debugging, resulting in K * K debugged proofs. For Accuracy@1, we use greedy decoding, which only generate 1 output, and for Accuracy@10, we sample 10 outputs with a temperature of 0.7. Then, for each task in our benchmarks, Accuracy@K equals one if at least one proofs/debugged proofs is verified by Verus.
@@ 4.2 Evaluation Results
4.2.1 Main Results
In Table 1, we compare the accuracy of SAFE with our baseline approaches. Due to the inherent complexity of this task, there is no pre-existing fine-tuned model to serve as a baseline. We employ GPT-4o, DeepSeekCoder-33B-Instruct model (Guo et al., 2024) and Llama3.1-8B-Instruct model (Dubey et al., 2024) with basic prompts same as what we use for the fine-tuned models as our baselines. To offer some advantages to the baselines, we also feed four examples for in-context learning in these baselines, which we do not use for SAFE fine-tuned models.
@@ 4.2.1 Main Results
Table 1: Accuracy of SAFE and baselines. SAFE0 is GPT-4o with a long prompt that is used to bootstrap SAFE; SAFE is LLaMa3.1 or DeepSeekCoder with three rounds of finetuning and a simple prompt; SAFE+ allows a round of self-debugging beyond SAFE’s proof generation.
@@
Raw SAFE0 Raw SAFE SAFE+ Raw SAFE SAFE+
VerusBench @1 11.51 25.90 3.60 46.04 52.52 9.35 43.17 49.64
@10 24.46 41.01 11.51 53.96 64.03 17.27 53.96 70.50
@100 - - - 55.40 67.63 - 59.71 79.14
CodeNet-Test @1 0.28 2.86 0.00 44.32 48.50 0.21 43.83 48.43
@@
Self- Evolving Metric Proof Generation Self-Debugging
@@ 4.2.2 Benefits from Self-Evolving
In Table 2, we use the performance on VerusBench to show how our self-evolving framework improves the capability of models. In this set of experiments, we use DeepSeekCoder-33B-Instruct for data synthesis and fine-tuning; Round 1, 2, 3 represent three models fine tuned from DeepSeekCoder-33B-Instruct. Table 2 illustrates that for every subset of VerusBench, the best Accuracy@1 and Accuracy@10 scores are achieved at the last two rounds of model evolution, which indicts the effectiveness of self-evolution. Besides, the round 1 model with a simple prompt already outperforms GPT-4o with a much more sophisticated prompt for all but the “Tutorial” subset. Round 3 model outperforms Round 2 model in terms of Accuracy@10 for all subsets of tasks, with or without self-debugging. Meanwhile, the improvement from Round 2 to Round 3 is much smaller than the improvement from GPT-4o to Round 1, and from Round 1 to Round 2, which justifies our decision of stopping the self-evolution after three rounds. We further report paired t-test results for Table 2 in Table 6 in Appendix.
Finally, the Round 3 model greatly outperforms GPT-4o for all three sub-sets of VerusBench, even though GPT-4o uses a much more sophisticated prompt: GPT-4o was most effective for the SV subset with a 39.47% Accuracy@1, while Round 3 model improves that metric to 78.95% (81.58% with self-debugging); GPT-4o was the least effective for the MBPP subset with a 19.23% Accuracy@1, which is improved by Round 3 model to 30.77% (41.03% with self-debugging). This trend bolds well for SAFE to work with different datasets and different bootstrapping methods in the future.
@@ 4.2.3 Improvement of Self-debugging
Table 1 demonstrates that SAFE+, which allows LLMs to do self-debugging, substantially improves the accuracy compared to direct generation, indicting the effectiveness of the self-debugging mechanism. In Table 2, self-debugging mechanism shows consistent improvements in all the three components of VerusBench. As we show in Listing 2 of the Appendix, Verus provides error messages for any incorrect proof about which proof annotation, which part of the specification, or which implicit proof target (e.g., no overflow for every arithmetic expression) cannot be verified, which can help repair the proof for experienced human users and well trained models.
Decoding Strategies for Self-Debugging. From Table 1, we can see the enhancements of self-debugging become more significant with the increase in the number of sampled outputs. To further evaluate how decoding strategies affect the performance of self-debugging in VerusBench, we conduct four settings by combining greedy and sampling decoding during the generation and debugging phases, and run two more self-debugging rounds with greedy decoding. Figure 2 shows that more rounds of self-debugging hardly improve the accuracy, with the improvement less than 1%. Besides, sampling decoding strategy is more useful for generation phase rather than for the self-debugging process. Notably, the “sampling+greedy” (generate multiple proofs and for each incorrect proof generate one debugged proof) strategy outperforms the “greedy+sampling” (generate one proof and generate multiple debugged proofs) setting.
@@
Figure 2: SAFE’s accuracy with different self-debugging, sampling decoding strategy

### v2

@@ 3.1 Step 1: Generating Verus-compatible Code
To address this compatibility issue, we employ GPT-4o as a code translator, effectively substituting Verus-incompatible Rust code snippets with Verus-compatible alternatives, such as converting an iterator-based implementation into a while-loop based implementation. Specifically, we prompt GPT-4o with all the unsupported features sourced from Verus official documentation. All the converted code that can pass the grammar check of Verus compiler is collected for next steps. This step is performed only once in our self-evolving framework.
@@ 3.2 Step 2: Self-evolving Specification Synthesis
In this step, LLMs are tasked with generating preconditions and postconditions for a Rust function based on the function’s implementation and docstring. Different from prior work in specification generation (Flanagan & Leino, 2001; Ma et al., 2024), SAFE needs a synthesized specification as input for its self-evolving proof-synthesis framework, which raises unique requirements on its specification evaluation criteria (i.e., which specification to keep or discard) and the mechanism to evaluate the criteria.
@@
In terms of criteria, a perfect specification $S$ should be correct (i.e., any correct implementation should be accepted by $S$) and complete (i.e., any incorrect code implementation should be rejected by $S$). In SAFE, getting perfect specification is not the goal. SAFE should discard incorrect specifications, which can never be proved and hence will waste training cycles in the next step. However, SAFE is fine with incomplete specifications, as the usage of specifications in SAFE is to stimulate proof synthesis not to judge code correctness. In fact, SAFE needs incomplete specifications, as complete specifications are likely too challenging to prove at the early stage of the proof-synthesis model’s evolution. Of course, specifications that are too incomplete should be discarded, because trivial specifications can be proved without any proof annotations and hence offer no usage for proof-synthesis training (e.g., 0 <= r for the binary search function).
@@
In terms of the mechanism to evaluate the preceding criteria, previous work relies on running many test cases (Endres et al., 2024), formally verifying the consistency between specifications and code (Ma et al., 2024), or user inspection (Lahiri et al., 2022). None of these mechanisms suit SAFE: user inspection or running many test cases would severely slow down the self-evolving framework; formal verification also does not work when we do not have a capable model to automatically synthesize proofs.
With all these constraints, SAFE chooses to leverage a recently proposed technique (Lahiri, 2024) to measure the Correctness and Completeness of specifications. We use Verus to symbolically evaluate what percentage of test cases $\mathbf{T}$ provided by the original dataset can pass a given specification $S$ (i.e., the Correctness score), and what percentage of test cases $\mathbf{T^{\prime}}$ mutated from $\mathbf{T}$ can be rejected by $S$ (i.e., the Completeness score). For a given method $m(x):y$, with $x$ being the input parameter and $y$ being the output, we denote the input/output-value pair of any test case in a test suite $\mathbf{T}$ as $(i,o)$ and generate a corresponding mutated test case by changing $o$ into a randomly different value $o^{\prime}$. The two metrics for a specification $S$ are formally defined as following, with $S(i,o)$ being a boolean value representing whether the input-output value pair $(i,o)$ satisfies the specification $S$ or not:
CorrectnessS: $\frac{\mid\{\ (i,o)\ |\ S(i,o)\ \ and\ \ \ (i,o)\in\mathbf{T}\}\mid}{\mid\mathbf{T}\mid}$ (1)
CompletenessS: $\frac{\mid\{\ (i,o^{\prime})\ |\ not\ S(i,o^{\prime})\ \ and\ \ \ (i,o^{\prime})\in\mathbf{T^{\prime}}\}\mid}{\mid\mathbf{T^{\prime}}\mid}$
@@
We can now set a threshold that fits the need of SAFE: we keep all the specifications with $\geq$80Correctness score and $\geq$60of Completeness score. This way, SAFE can obtain many imperfect but useful specifications for its proof synthesis. We do not set the Correctness threshold to be 100%, because we use GPT to generate test cases when few test cases are provided by the dataset, and we cannot guarantee perfect correctness of these test cases. Since Verus can symbolically evaluate $S(i,o)$ in the preceding formula without checking or running the Rust function, the scoring is fast.
@@
For each function, we keep up to three synthesized specifications, which not only enrich the specification training dataset but also increase the chance of synthesizing valid proofs later. Our self-evolving process for specification stops when there is no significant increase in the total number of accepted function–specification pairs, at which time all these pairs move on to become the input for the next step.
@@ 3.3 Step 3: Self-evolving Proof Synthesis
Now that we have tens of thousands of function-specification pairs, we can start the self-evolving procedure of proof synthesis. Compared with specification synthesis, deciding whether to accept or reject a synthesized item is much more straightforward—only proof that Verus can use to verify the Rust function satisfies its specification is accepted; if any verification error is raised, the proof is discarded [^1] If a Rust function is in fact buggy or a synthesized specification is incorrect, no proof will be accepted.. However, since proof synthesis is much more difficult than specification synthesis, we need to pay special attention to bootstrap the self-evolving procedure and to keep it moving forward.
For bootstrapping, a simple prompt to GPT-4o would not work, as it is very difficult for GPT-4o to synthesize all the needed loop invariants without any incorrect ones in between, not to mention more complicated proof annotations like proof blocks and lemma functions. Therefore, we write a detailed prompt for GPT-4o, explaining the principles and tricks of writing Verus proofs. By doing so, we finally manage to make GPT-4o synthesize proof annotations that allow Verus to prove more than one thousand Rust functions at the cost of one whole month of non-stop GPT-4o invocations. Fortunately, doing so is sufficient for bootstrap. For the remainder of the self-evolving procedure, GPT-4o is not used any more. Instead, we use an open-source LLM much smaller than GPT-4o to efficiently synthesize data and improve its capability in a self-evolving manner as in Algorithm 1. Specifically, we fine-tune our open-source LLM on two tasks, proof generation and self-debugging.
Task #1: Proof generation. The synthesized proofs and specifications are directly used for training this task. Taking a code snippet with specifications as input $S$, LLM $M$ is tasked with generating a proof $Y$ (consisting of tokens $y_{0},\dots,y_{N}$) that can make the code proved by Verus. We train this task with a sequence-to-sequence objective where
$\vspace{-0.2cm}\mathcal{L}_{Gen}(\theta)=-\sum_{i=0}^{N}\log P_{\theta}(y_{i}\mid S,y_{t<i})\vspace{-0.2cm}$ (2)
@@
$\mathcal{L}_{Debug}(\theta)=-\sum_{i=0}^{N}\log P_{\theta}(y_{i}\mid Y_{X},\mbox{{{Error}}}_{Y_{X}},y_{t<i})$ (3)
It is important to train the model to “debug” an imperfect proof. In many cases, the proof annotations generated by the model contain only small errors or miss one line of annotation. Without the capability of self-debugging, the model would start from scratch and fail to synthesize a perfect proof even after many attempts. In SAFE, we conduct joint training for the proof generation task and the self-debugging task, empowering the LLM to both generate a proof from scratch and repair an existing proof based on verification error messages for a Rust function and its specification.
@@ 4.1 Dataset
As an early exploration of automated formal verification, in this paper, we focus solely on the Rust code of algorithm types at the function level, as our source of data. We employ the MBPP dataset (Austin et al., 2021) (only training split for data synthesis) and the CodeNet dataset (Puri et al., 2021) as our data sources. These two datasets contain small programs written in different programming languages, each of which is an intended solution to a coding problem described in natural language and is associated with three test cases on average. We translate the Python programs in MBPP and extract the Rust programs in CodeNet. In total, we collect 45,395 Rust single-function programs. Of these programs, 21,398 have been successfully transformed into Rust code that is compatible with Verus by GPT-4o. We conduct the specification synthesis first, and after two rounds of self-evolution, we obtain 19,017 high-quality specifications. Then we run three rounds of self-evolving proof synthesis, ending with 9,706 verified programs, and 10,486 self-debugging data pairs. To the best of our knowledge, this dataset represents the most extensive synthetic dataset created for Verus code to date.
4.2 Benchmark and Metrics
@@
We assess a model’s proficiency in generating proofs in two benchmarks, the human-written benchmark VerusBench and the synthetic benchmark CodeNet-Test. VerusBench is a human-written benchmark dataset with human-written Rust code and corresponding Verus specifications. It contains 139 code files in total. 23 of them are algorithmic programs from Verus tutorials. 38 of them come from the dataset of SV-COMP-2021 (Beyer, 2021) (in short, SV), which is a contest focused on program verification for C and Java languages; we utilize 38 Verus-translated tasks provided by Yao et al. (2023). The remaining 78 tasks come from MBPP-DFY-153 (Misu et al., 2024), a Dafny version of the MBPP test set. We have translated 78 of them that are compatible with Verus. CodeNet-Test is an expansive benchmark, 10x larger than VerusBench, crafted by LLM and encompassing 1,435 diverse tasks. Given the substantial human effort required to write Verus code and specifications, it’s impractical to create a large test suite by human. To comprehensively measure the model capability of generating proofs, we split a subset of CodeNet for testing, ensuring that it is not utilized in our data synthesis process. We leverage our specification generator to craft specifications for this subset and apply the same scoring mechanism to preserve reasonably well specifications.
@@
We measure the capability of proof generation by Accuracy@K. Specifically, we use any model under evaluation to sample K proofs during proof generation. For Accuracy@1, we use greedy decoding, which generates only 1 output, and for Accuracy@10, we sample 10 outputs with a temperature of 0.7. Then, for each task in our benchmark set, Accuracy@K equals one if at least one proof/debugged proof is verified by Verus. Note that, when the self-debugging feature of SAFE is used (i.e., SAFE+), the proof generation includes two rounds: at the first round, $K$ initial proofs are sampled; at the second round, if none of the initial proofs are correct, $K$ proofs are generated by self-debugging each initial proof, resulting in K * K debugged proofs.
@@ 4.3 Evaluation Results
In Table 1, we compare the accuracy of SAFE with our baseline approaches. Due to the inherent complexity of this task, there is no pre-existing fine-tuned model to serve as a baseline. We employ GPT-4o, DeepSeekCoder-33B-Instruct model (Guo et al., 2024), and Llama3.1-8B-Instruct model (Dubey et al., 2024) with basic prompts the same as what we use for the fine-tuned models as our baselines. To offer some advantages to the baselines, we also feed four examples for in-context learning in these baselines, which we do not use for SAFE fine-tuned models.
@@
Table 1 also shows the efficacy of self-debugging (i.e., SAFE+). For example, after one initial proof is sampled, applying self-debugging on this proof to produce the second sample (Accuracy@2 for SAFE+) consistently outperforms simply producing a second sample of proof without debugging (Accuracy@2 for SAFE) for both LLaMa and DeepSeekCoder on both VerusBench and CodeNet-Test, as shown in Table 1. SAFE+ with DeepSeekCoder achieves the best performance of 70.50% in Table 1 when 10$\times$10 debugged proofs are produced for 10 initial proofs for each proof task.
Overall, by leveraging our self-evolving framework, SAFE can substantially improve the capability of open-source model and effectively generate proofs for Verus programs.
Table 1: Accuracy of SAFE and baselines. Prompt is GPT-4o with a long prompt used to bootstrap SAFE; column SAFE is LLaMa3.1 or DeepSeekCoder with three rounds of finetuning and a simple prompt. For SAFE+, Accuracy@2 means getting one initial proof sample and one debugging sample; Accuracy@100 means getting $10$ initial proofs and then $10\times 10$ debugging samples.
@@
Raw Prompt Raw SAFE SAFE+ Raw SAFE SAFE+
VerusBench @1 11.51 25.90 3.60 46.04 - 9.35 43.17 -
@2 14.39 30.93 5.76 48.20 52.52 10.79 46.76 49.64
@10 24.46 41.01 11.51 53.96 - 17.27 53.96 -
@100 43.88 46.76 28.78 55.40 64.03 32.37 59.71 70.50
CodeNet-Test @1 0.28 2.86 0.00 44.32 - 0.21 43.83 -
@2 0.70 3.41 0.03 45.34 48.50 0.55 44.74 48.43
@@
Self- Evolving Metric Proof Generation Self-Debugging ($K+K*K$)
@@ 4.3.1 Benefits from Self-Evolving
In Table 2, we use the performance on VerusBench to show how our self-evolving framework improves the capability of models. In this set of experiments, we use DeepSeekCoder-33B-Instruct for data synthesis and fine-tuning; Rounds 1, 2, 3 represent three models fine tuned from DeepSeekCoder-33B-Instruct. Table 2 illustrates that for every subset of VerusBench, the best Accuracy@1 and Accuracy@10 scores are achieved at the last two rounds of model evolution, indicating the effectiveness of self-evolution. Besides, Round 1 model with a simple prompt already outperforms GPT-4o with a much more sophisticated prompt for all but the “Tutorial” subset. Round 3 model outperforms Round 2 model in terms of Accuracy@10 for all subsets of tasks, with or without self-debugging. Meanwhile, the improvement from Round 2 to Round 3 is much smaller than the improvement from GPT-4o to Round 1, and from Round 1 to Round 2, justifying our decision of stopping the self-evolution after three rounds. We further report paired t-test results for Table 2 in Table 6 in Appendix.
Finally, Round 3 model greatly outperforms GPT-4o for all three sub-sets of VerusBench, even though GPT-4o uses a much more sophisticated prompt: GPT-4o is most effective for the SV subset with a 39.47% Accuracy@1, while Round 3 model improves that metric to 78.95% (81.58% with self-debugging); GPT-4o is the least effective for the MBPP subset with 19.23% Accuracy@1, which is improved by Round 3 model to 30.77% (41.03% with self-debugging). This trend bolds well for SAFE to work with different datasets and different bootstrapping approaches in the future.
@@ 4.3.2 Improvement of Self-debugging
Table 1 demonstrates that SAFE+, which allows LLMs to do self-debugging, substantially improves the accuracy compared to direct generation, indicating the effectiveness of the self-debugging mechanism. As we show in Listing 2 of the Appendix, Verus provides error messages for any incorrect proof about which proof annotation, which part of the specification, or which implicit proof target (e.g., no overflow for every arithmetic expression) cannot be verified. These error messages can help repair the proof for experienced human users and well trained models.
Decoding Strategies for Self-Debugging. From Table 1, we can see that the enhancements of self-debugging become more significant with the increase in the number of sampled outputs. To further evaluate how decoding strategies affect the performance of self-debugging in VerusBench, we conduct four settings by combining greedy and sampling decoding during the generation and debugging phases, and run two more self-debugging rounds with greedy decoding. Figure 2 shows that more rounds of self-debugging hardly improve the accuracy, with the improvement less than 1%. Besides, the sampling decoding strategy is more useful for the generation phase rather than for the self-debugging phase. Notably, the “sampling+greedy” (generating multiple proofs and for each incorrect proof generate one debugged proof) strategy outperforms the “greedy+sampling” (generating one proof and generate multiple debugged proofs) setting. Existing work (Chen et al., 2023; Olausson et al., 2023) has shown similar results.
@@
Figure 2: SAFE’s accuracy with different self-debugging, sampling decoding strategies